find five numbers in A.P whose sum is 12/2 and their ratio of first to last is 2:3?
Answer by richwmiller(17219)

12/5 = (a + z),
a/z=2/3
a = 24/25, z = 36/25
a = 24/25, z = 36/25
36/25 = 24/25 + 4d
12/25=4d
3/25=d
24/25,27/25,30/25,33/25,36/25
an = 24/25 + 3/25(n - 1)